OpenStudy (anonymous):

how many real solutions does the equation 0.2x^5-2x^3+1.8x+k=0. have when k=0. a. how many real solutions when it k=1. b.is there a value for k for which the equation has just one real solution? b. is there a value for k for which the equation has no real solution?

OpenStudy (anonymous):

When k =0: \[0.2x^{5}-2x^{3}+1.8x=0\]\[x(0.2x^{4} - 2x^{2}+1.8)=0\]\[0.2x^{4}-2x^{2}+1.8=0\]Suppose y = x^2. Now, \[0.2y^{2}-2y+1.8=0\] Comparing it with ay^2+by+c=0: a=0.2 b=-2 c=1.8 \[y = \frac{ -b \pm \sqrt{b^{2}-4ac} }{ 2a }\]\[y = \frac{ 2\pm \sqrt{4-(4\times0.2\times1.8)} }{ 2\times0.2 }\]\[y = \frac{ 2\pm \sqrt{2.56} }{ 0.4 }\]Now using (+ve) and (-ve) sign find out the soultion for y and then put the results here:\[y = x^{2}\]You will get two values of x. Follow the same process for next one.

OpenStudy (anonymous):

When k=1 you can't aggregate the x... @curiousshubham

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

my head is burning...........

OpenStudy (anonymous):

y

OpenStudy (anonymous):

Today I read a lot.... and em feeling sleepy..

OpenStudy (anonymous):

good night

OpenStudy (anonymous):

you too.......

OpenStudy (anonymous):

n u?

OpenStudy (anonymous):

I feel like I'm trying to create a new math formula XD

OpenStudy (anonymous):

@decripter37 whats that?

OpenStudy (anonymous):

I following this path: ax^5+bx^3+cx=k => x(ax^4+bx^2+c)=k => x=k/y, ax^4+bx^2+c=y and I'm trying to find y :)

OpenStudy (anonymous):

keep it up...

OpenStudy (anonymous):

You can also draw the function, it's easy you just have to find max, min with derivatives, and after see how many times the function pass trough the 1

OpenStudy (anonymous):

ok thnks